# Space Group Informed Transformer for Crystalline Materials Generation

### v2

@@ II.3 Sampling
Besides autoregressive sampling, one can also perform Markov chain Monte Carlo (MCMC) sampling based on the likelihood Eq. (1) of the CrystalFormer. MCMC sampling of crystals is a way to walk through the crystalline materials starting from an existing crystal structure. At each step of the random walk, one proposes an element substitution, atom position shift, or lattice deformation to change the crystal from $\boldsymbol{\mathcal{C}}$ to $\boldsymbol{\mathcal{C}}^{\prime}$, then accepts or rejects the proposal according to the model probability according to the Metropolis acceptance rate $\min\left[1,\frac{p(\boldsymbol{\mathcal{C}}^{\prime}|g)}{p(\boldsymbol{\mathcal{C}}|g)}\right]$. MCMC sampling is particularly useful for incorporating additional constraints or guidance in the sampling procedure. Moreover, during the burn-in phase of such MCMC sampling, the generated samples will be similar to the starting material, which may be a desired feature in certain cases.
III CrystalFormer learns chemical intuition by compressing materials database
@@
Nature tends to favor symmetrical crystal structures. Crystallographic space groups capture and quantify this inductive bias of nature, thereby significantly simplifying the spaces of crystal materials. In light of the space group symmetries, crystals also have an unexpected yet natural sequential and discrete representation, which derives from two tables in nature: the periodic table of elements determined by quantum mechanics and the table of Wyckoff positions of the 230 space groups determined by group theory. To construct a certain crystal, we only need to select atoms from the periodic table and place them sequentially into the Wyckoff positions in the unit cell. In this crystal language, the “word order” is determined by the alphabetical order of Wyckoff letters, the “grammar” corresponds to the solid-state chemistry rules, and the “synonyms” represent interchangeable elements (Sec. III.1), the “sentence length” correspond to atom number in the unit cell (Sec. III.2), and the “idioms” correspond to common chemical coordination (Sec. III.3).
@@
CrystalFormer employs an autoregressive transformer to learn the crystal language, thereby exploring yet-to-be-discovered crystalline materials. It compresses and internalizes the crystal materials database, expressing solid-state chemical knowledge through neural network parameters; reflecting the associative ability of material space through neural network activations; and describing chemical intuition through the model probability (Sec. III.4). Similar to generative models used for generating text, images, and videos, CrystalFormer can directly generate “realistic” crystal materials. However, rather than worrying about the fake contents of AI-generated media, these AI-generated crystal materials could potentially be synthesized and be useful to human civilization.
@@
Next, we will inspect the learned features and sample statistics of the model to build up an understanding of the CrystalFormer. We carry out inspectations for a few selected space groups. The findings are neverthelss general. These findings not only provide confidence a bit of understandings of the model, but also direct us to the suitable applications of CrystalFormer.
III.1 Atom embeddings and chemical similarity
@@
Figure 3 visualizes the cosine similarity of the learned vector embedding of the chemical species. Red colors in the figure indicate similar chemical species identified by the model. One sees the chemical similarity within groups of elements show up as off-diagonal red stripes. Moreover, there are visible clusters for Lanthanide elements (La-Lu). The plot also suggests the similarity between the lanthanides and other rare-earth elements (Y and Sc). The features shown in Fig. 3 are strikingly similar to the similarity map constructed purposely based on substitution pattern [46, 47] which was later used for substitution-based material discoveries [48, 5]. In the context of language modeling, the chemical similarities correspond to synonyms of chemical species tokens. Having the ability to learn them from data  [46, 47, 49, 50, 51, 52, 53, 19] is an encouraging signal that the model is picking up atomic physics to be able to generate reasonable crystal structures.
III.2 Atom number distributions
@@
The number of atoms corresponds to the length of non-padding atoms in the sequential representation, which is captured well by CrystalFormer. Figure 4 presents the histogram of the total number of atoms in the conventional unit cell for several space groups. One sees a nice agreement between the atom number distribution in the test dataset and the generated samples. In addition, it appears that space group $g$ is the key latent variable that decomposes the multi-modal atom number distribution of crystals. This is understandable because the number of atoms is determined by the sum of the multiplicities of occupied Wyckoff positions. Therefore, the space group symmetry is a key control variable for the atom number distribution. Incorporating Wyckoff positions information into the CrystalFormer model architecture removes the necessity of querying the training data to find out the number of atoms for a targeted space group [16].
@@
Figure 4: The histogram for total number of atoms in the unit cell for several space groups in the test dataset and in the generated samples.
@@
Figure 5: The heat map for Wyckoff positions and atom species of (a) the test dataset and (b) generated samples for the $Fm\bar{3}m$ space group (No. 225). It is an analog of bigram frequency statistics of language modeling, which shows where atoms tend to occupy in the unit cell.
@@
Recently, Ref. [54] reports an abundance of inorganic compounds whose primitive unit cell contains a number of atoms that is a multiple of four. There are different ways to reason about the observed ”rule of four” depending on one’s view of how a crystal is formed. For example, one can often break inorganic solids into polyhedra as building blocks. Otherwise, Ref. [55] considers the most probable values of the number of atoms in a formula unit and the number of formula units per primitive cell. In line with the discussion here, the ”rule of four” is the combination of three factors 1) the distribution of crystalline materials among space groups [30]; 2) the distribution of atoms in Wyckoff positions [33] of a given space group; and 3) the multiplicities of Wyckoff positions and multiplicities of conventional versus cells. The first two are statistical rules determined by the inter-atomic interactions while the third one is a mathematical fact of space group theory. In the end, the point we want to make is that the design of CrystalFormer and its associated crystal representation allow it to learn the ”rule of four” and many other to-be-discovered ”rules”, which manifest themselves as marginal statistics of learn probability distribution. Most importantly, CrystalFormer will utilize these ”empirical rules” when generating novel yet reasonable crystal samples.
III.3 Wyckoff-Atom gram
@@
Figure 5 shows heat maps of Wyckoff positions and chemical species for the $Fm\bar{3}m$ space group (No. 225). The heat map is analogous to bigram frequency statistics in language modeling. In the present context, it reveals interesting solid-state chemistry knowledge related to where each atom tends to appear in a unit cell. First of all, one sees that most atoms occupy special Wyckoff positions (Wyckoff letters at the beginning of the alphabet) with higher site symmetries. The distribution of generated data is in agreement with test data and recent statistics [33]. Moreover, there are vertical blanks at the locations of inert elements (He, Ne, Ar…) as they are rare in crystalline materials. Lastly, one sees that oxygen and halogen elements (F, Cl, Br, I) appear quite often in the Wyckoff position ”24e”, which is a consequence of their electronegativities [30]. Overall, we see the CrystalFormer has learned these key motifs for generating crystalline materials. Nevertheless, one sees that several Wyckoff locations of the hydrogen are missing in the generated samples compared to the test dataset. We believe that is due to that the hydrogen element takes only about $0.4\%$ in the training data for the $Fm\bar{3}m$ space group. Collecting more data with better coverage of elements will be crucial to further boost the performance of the current model.
@@
III.4 Crystal likelihoods
Figure 6: The log-likelihood histogram for 1000 samples in the $Fm\bar{3}m$ (No. 225) space group and the test dataset. The insets visualize the crystal structure of a few generated samples. Rb2AgBiBr6 and LiMg2Cd are in the training dataset. ErAs is in the validation dataset. K2YbCeF6 and Ba6CoNi(IN3)2 are not in the MP-20 or Materials Project database.
CrystalFormer compresses chemistry knowledge stored in the material dataset into its parameters. In addition to generating crystal samples, CrystalFormer can also compute the likelihoods of crystals via Eq. (1). Therefore, it is possible to employ CrystalFormer in the Monte Carlo search of crystal structures besides sampling materials directly as a generative model.
Figure 6 shows the agreement of the likelihoods of generated samples and samples in the test dataset. We also visualize structures of a few generated samples which are deemed to be very likely, typical, and unlikely according to their likelihood values. We have checked that likelihood is related to the energy of the crystal by locally perturbing the fractional coordinates and lattice parameters. However, we did not observe a correlation between the likelihood of these crystals and their energies on a global scale. We envision the landscape of likelihoods is much less rough compared to the potential energy surface of crystalline materials. Intuitively, it means that the Crystalformer compresses the materials space into a more compact space without many holes that correspond to infeasible high energy states. Therefore, likelihood-based exploration of the crystal space discussed in Sec. II.3 can be more efficient compared to traditional sampling approaches based on the physical energy functions.
IV Applications
@@
IV.1 Symmetry-conditioned random structure initialization
@@
Figure 7: (a)-(g) Average energy difference versus relaxation steps for seven representative space groups. (h) The histograms of energy above the convex hull for relaxed crystal structures. The dashed line indicates the criterion for selecting candidates for stable materials listed in Appendex C since materials with $E_{\mathrm{hull}}<0.1$ eV/atom are usually metastable and have the potential to be synthesized [74].
Space group Crystal system Converged structures $\uparrow$ Retain symmetry $\uparrow$ RMSD 1 ($\AA$) $\downarrow$ $E_{\mathrm{hull}}$ 1 (eV/atom) $\downarrow$
@@
$P\bar{1}$ (No. 2) Triclinic 46 67 45 67 1.929 2.038 1.034 0.913
$C2/m$ (No. 12) Monoclinic 55 72 53 67 1.846 2.235 1.233 1.660
$Pnma$ (No. 62) Orthorhombic 77 83 76 66 1.086 1.940 0.313 1.633
$I4/mmm$ (No. 139) Tetragonal 91 81 88 63 0.405 1.160 0.240 1.100
$R\bar{3}m$ (No. 166) Trigonal 83 74 80 71 1.028 2.183 0.352 2.489
$P6_{3}/mmc$ (No. 194) Hexagonal 97 77 96 60 0.355 2.129 0.324 4.100
$Fm\bar{3}m$ (No. 225) Cubic 98 96 95 92 0.076 0.518 0.214 0.483
Average 78.1 78.6 76.1 69.4 0.961 1.743 0.530 1.769
@@
Figure 7(a)-(g) shows the average energy difference to the energy of final structures versus DFT relaxation steps. We neglected the structures whose energy changes and energy change intervals per step during relaxations exceeded 10 eV/atom to eliminate the impact of erroneous steps. One sees that CrystalFormer samples generally reach lower energies in fewer relaxation steps. This is especially true for space groups with higher symmetries. The ability to initialize diverse and high-quality crystal structures enables one to discover more stable materials faster. Figure 7(h) shows the histogram of energy above the convex hull constructed by the Materials Project database. The dashed line denotes the criterion $E_{\mathrm{hull}}<0.1$ eV/atom [74] for selecting stable materials. Among these candidates, we found 34 and 12 relaxed structures with CrystalFormer and PyXtal initializations that are not contained in the MP-20 dataset. We summarize them in Table LABEL:table:samples and Table LABEL:table:pyxtal_samples of appendix C.
Table LABEL:table:700-stat lists detailed statistics of structure-relaxed samples in seven representative space groups. Overall, we find that the CrystalFormer generated structures are of higher quality, especially for those space groups with higher symmetry. This observation is supported by the fact that the DFT relaxation often retains the space group symmetry. The average root mean squared displacement (RMSD) computed for these converged structures is 0.961 $\AA$ and 1.743 $\AA$ for CrystalFormer and PyXtal initialization, respectively. In addition, the average energy above the convex hull also confirms the samples generated by CrystalFormer are indeed much closer to the DFT local minimum than PyXtal initialization.
IV.2 Structure-conditioned element substitution
@@
Many crystal structures can be traced back to a few simple, highly symmetrical types. Numerous crystals share the same structural prototype but differ in composition, such as perovskite (ABX3), spinel (AB2X4), fluorite (AX2), and so on. Figure 8(a) shows double perovskite crystal structures A2BB’X6 which belong to the $Fm\bar{3}m$ (No. 225) space group. There are hundreds of known double perovskites with significant interests in their semiconducting, ferroelectric, thermoelectric, and superconducting properties [78]. Finding more stable materials with this structure prototype using brute force enumeration and high-throughput calculation is a computationally demanding task [79]. We will generate new double perovskites with CrystalFormer and demonstrate its advantage of over standard element substitution methods.
Figure 8(a) shows the string representation of double perovskites. To generate candidates of double perovskites, we use CrystalFormer to carry out string in-filling tasks. Since the autoregressive sampling of the atoms is insufficient to take into account non-causal information in the sequence, we employ MCMC to sweep through the sequence and update chemical species and fractional coordinates [80]. The acceptance rate for these MCMC updates makes use of the marginalized probability for elements and fractional coordinates as the lattice parameter that appears at the end of the sequence can be integrated. Only after the MCMC sampling has been thermalized, we sample the lattice parameters autoregressive to account for the adjustment of the unit cell for given atoms and occupations. We use CrystalFormer to generate 100 candidates as the initial DFT relaxation.
@@
The RMSD computed for the DFT-relexed structures is 0.211 $\AA$ and 0.342 $\AA$ for CrystalFormer and ionic substitution [46], respectively. This observation confirms that the structures generated by CrystalFormer are closer to the DFT local minimum than ionic substitution [46]. Moreover, Figure 8(b) shows the histogram of energy above the convex hull of the Materials Project database. Overall, CrystalFormer and ionic substitution [46] found 9 and 3 double perovskites with $E_{\mathrm{hull}}<0.1$eV/atom which are not contained in the MP-20 dataset, details in appendix D. The superior performance of CrystalFormer-guided MCMC is understandable since its likelihood takes into account the context of space group and atomic environment rather than marginal two-body correlation [46] in ionic substitution. The ionic substitution approach also shows two limitations in practical applications. First, some of the ions in the compound can not be substituted as they are missing in the probability table. Second, the approach relies on the calculability of the elements’ valence states.
@@
Figure 8: (a) Double perovskites crystal structure. The crystal string representation of double perovskites with blank spaces for chemical elements and the x-coordinate of the atom resides in the ’e’ position. CrystalFormer generates crystals with double perovskite structures via sequence infilling. (b) The histograms of energy above the convex hull for the relaxed crystal structures. The dashed line indicates the criterion for selecting candidates for the stable materials.
IV.3 Plug-and-play materials design
Figure 9: (a) The histogram of band gap and formation energy for crystal samples generated in the $Fm\bar{3}m$ space group (No. 225). The dashed red lines in the plane indicate target values. The marginals on the side show the shift of the property distributions with respect to the unconditionally generated samples. Note that we scale the 3d histograms for better visualization. (b) The likelihoods of conditioned generated samples compared to the unconditional samples.

### v3

@@ II.1.3 Sampling
Besides autoregressive sampling, one can also perform Markov chain Monte Carlo (MCMC) sampling based on the likelihood Eq. (1) of the CrystalFormer. MCMC sampling can walk through the crystalline materials space starting from an existing crystal structure. At each step of the random walk, one proposes a configuratoin update in terms of element substitution, atom position shift, or lattice deformation to change the crystal from $\boldsymbol{\mathcal{C}}$ to $\boldsymbol{\mathcal{C}}^{\prime}$, then accepts or rejects the proposal according to the model probability following the Metropolis acceptance rule $\min\left[1,\frac{p(\boldsymbol{\mathcal{C}}^{\prime}|g)}{p(\boldsymbol{\mathcal{C}}|g)}\right]$. MCMC sampling is particularly useful for incorporating additional constraints or guidance in the sampling procedure. Moreover, during the burn-in phase of such MCMC sampling, the generated samples will be similar to the starting material, which may be a desired feature in certain cases.
II.2 CrystalFormer learns chemical intuition by compressing materials database
@@
Nature favors symmetrical crystal structures. Crystallographic space groups quantify this inductive bias of nature, thereby significantly simplifying the spaces of crystal materials. In light of the space group symmetries, crystals also have an unexpected yet natural sequential and discrete representation, which derives from two tables in nature: the periodic table of elements determined by quantum mechanics and the table of Wyckoff positions of the 230 space groups determined by group theory. To construct a certain crystal, we only need to select atoms from the periodic table and place them sequentially into the Wyckoff positions in the unit cell. In this crystal language, the “word order” is determined by the alphabetical order of Wyckoff letters, the “grammar” corresponds to the solid-state chemistry rules, and the “synonyms” represent interchangeable elements (Sec. II.2.1), the “sentence length” correspond to atom number in the unit cell (Sec. II.2.2), and the “idioms” correspond to common chemical coordination (Sec. II.2.3).
@@
CrystalFormer employs an autoregressive transformer to learn the crystal language, thereby exploring yet-to-be-discovered crystalline materials. It compresses and internalizes the crystal materials database, expressing solid-state chemical knowledge through neural network parameters; reflecting the associative ability of material space through neural network activations; and describing chemical intuition through the model probability (Sec. II.2.4). Similar to generative models used for generating text, images, and videos, CrystalFormer can directly generate “realistic” crystal materials. However, rather than worrying about the fake contents of AI-generated media, these AI-generated crystal materials could potentially be synthesized and be useful to human civilization.
@@
Next, we will inspect the learned features and sample statistics of the model to build up an understanding of the CrystalFormer. We carry out inspectations for a few selected space groups. The findings are neverthelss general. These findings provide understandings and confidence of the model, therefore direct us to the suitable applications of CrystalFormer.
II.2.1 Atom embeddings and chemical similarity
@@
Fig. 3 visualizes the cosine similarity of the learned vector embedding of the chemical species. Red colors in the figure indicate similar chemical species identified by the model. One sees the chemical similarity within groups of elements show up as off-diagonal red stripes. Moreover, there are visible clusters for Lanthanide elements (La-Lu). The plot also suggests the similarity between the lanthanides and other rare-earth elements (Y and Sc). The features shown in Fig. 3 are strikingly similar to the similarity map constructed purposely based on substitution pattern [45, 46] which was later used for substitution-based material discoveries [47, 5]. In the context of language modeling, the chemical similarities correspond to synonyms of chemical species tokens. Having the ability to learn chemical similarities from data  [45, 46, 48, 49, 50, 51, 52, 19] is an encouraging signal that the model is picking up atomic physics for generating reasonable crystal structures with maximum likelihood based training.
II.2.2 Atom number distributions
@@
The number of atoms in the unit cell corresponds to the length of non-padding atoms in the crystal string representation, which is captured well by CrystalFormer. Fig. 4 presents the histogram of the total number of atoms in the conventional unit cell for several space groups. One sees a nice agreement between the atom number distribution in the test dataset and the generated samples. In addition, it appears that space group $g$ is the key latent variable that decomposes the multi-modal atom number distribution of crystals. This is understandable because the number of atoms is determined by the sum of the multiplicities of occupied Wyckoff positions. Therefore, the space group symmetry is a key control variable for the atom number distribution. Incorporating Wyckoff positions information into the CrystalFormer model architecture removes the necessity of querying the training data to find out the number of atoms for a targeted space group [16] during generation.
@@
Figure 4: The histogram for total number of atoms in the unit cell for (a) $P\bar{1}$ (b) PnmaPnma (c) $I4/mmm$ (d) $Fm\bar{3}m$ space groups in the test dataset and in the generated samples.
@@
Figure 5: The heat map for Wyckoff positions and atom species of (a) the test dataset and (b) generated samples for the $Fm\bar{3}m$ space group (No. 225). It is an analog of bigram frequency statistics of language modeling, which shows the prefereed Wyckoff positions of different chemical species.
@@
Recently, Ref. [53] reports an abundance of inorganic compounds whose primitive unit cell contains a number of atoms that is a multiple of four. There are different ways to reason about the observed “rule of four” depending on one’s view of how a crystal is formed. For example, one can often break inorganic solids into polyhedra as building blocks. Otherwise, Ref. [54] considers the most probable values of the number of atoms in a formula unit and the number of formula units per primitive cell. In line with the discussion here, the “rule of four” is the combination of three factors 1) the distribution of crystalline materials among space groups [30]; 2) the distribution of atoms in Wyckoff positions [33] of a given space group; and 3) the multiplicities of Wyckoff positions and multiplicities of conventional versus primitive cells. The first two are statistical rules determined by the inter-atomic interactions while the third one is a mathematical fact of space group theory. In the end, the point we want to make is that the design of CrystalFormer and its associated crystal representation allow it to learn the “rule of four” and many other to-be-discovered “rules”, which manifest themselves as marginal statistics of learned probability distribution. Most importantly, CrystalFormer will utilize these “empirical rules” when generating novel yet reasonable crystal samples.
II.2.3 Wyckoff-Atom gram
@@
Fig. 5 shows heat maps of Wyckoff positions and chemical species for the $Fm\bar{3}m$ space group (No. 225). The heat map is analogous to bigram frequency statistics in language modeling. In the present context, it reveals interesting solid-state chemistry knowledge related to where each atom tends to appear in a unit cell. First of all, one sees that most atoms occupy special Wyckoff positions (Wyckoff letters at the beginning of the alphabet) with higher site symmetries. The distribution of generated data is in agreement with test data and recent statistics [33]. Moreover, there are vertical blanks at the locations of inert elements (He, Ne, Ar…) as they are rare in crystalline materials. Lastly, one sees that oxygen and halogen elements (F, Cl, Br, I) appear quite often in the Wyckoff position “24e”, which means these high electronegative elements form polyhedra enviorment for other atoms [30]. Overall, we see the CrystalFormer has learned these key motifs for generating crystalline materials. On the other hand, one also observes that several Wyckoff locations of the hydrogen are missing in the generated samples compared to the test dataset. We believe that is due to that the hydrogen element takes only about $0.4\%$ in the training data for the $Fm\bar{3}m$ space group. Collecting more data with better coverage of elements will be crucial to further boost the performance of the current model.
@@
II.2.4 Crystal likelihoods
Figure 6: The histogram of log-likelihoods of 1000 samples in the $Fm\bar{3}m$ (No. 225) space group and the test dataset. The insets visualize the crystal structure of a few generated samples. Rb2AgBiBr6 and LiMg2Cd are in the training dataset. ErAs is in the validation dataset. K2YbCeF6 and Ba6CoNi(IN3)2 are not in the MP-20 or Materials Project database.
CrystalFormer compresses chemistry knowledge stored in the material dataset into its parameters. In addition to generating crystal samples, CrystalFormer can also compute the likelihoods of crystals via Eq. (1). Therefore, it is also possible to employ CrystalFormer in likelihood-based Monte Carlo search besides sampling crystals directly.
Fig. 6 shows the agreement of the likelihoods of generated samples and samples in the test dataset. We also visualize structures of a few generated samples which are deemed to be very likely, typical, and unlikely according to their likelihood values. We have checked that likelihood is related to the energy of the crystal by locally perturbing the fractional coordinates and lattice parameters. However, we did not observe a correlation between the likelihood of these crystals and their energies on a global scale. We envision the landscape of likelihoods is much less rough compared to the potential energy surface of crystalline materials. Intuitively, it means that the Crystalformer compresses the materials space into a more compact space without many holes that correspond to infeasible high energy states. Therefore, likelihood-based exploration of the crystal space discussed in Sec. II.1.3 can be more efficient compared to traditional sampling approaches based on the Boltzmann distribution based on physical energy functions.
III Results
@@
III.1 Symmetry-conditioned random structure initialization
@@
Figure 7: (a)–(g) Average energy difference versus relaxation steps for seven representative space groups. The insets show the distribution of ionic steps. (h) The histograms of energy above the convex hull for relaxed crystal structures. The dashed line indicates the criterion for selecting candidates for stable materials listed in Appendex C since materials with $E_{\mathrm{hull}}<0.1$ eV/atom are usually metastable and have the potential to be synthesized [73].
[…]
@@
$P\bar{1}$ (No. 2) Triclinic 46 67 45 67 1.181 1.259 1.034 0.913
$C2/m$ (No. 12) Monoclinic 55 72 53 67 1.051 1.227 1.233 1.660
PnmaPnma (No. 62) Orthorhombic 77 83 76 66 0.594 1.092 0.313 1.633
$I4/mmm$ (No. 139) Tetragonal 91 81 88 63 0.140 0.675 0.240 1.100
$R\bar{3}m$ (No. 166) Trigonal 83 74 80 71 0.294 0.919 0.352 2.489
$P6_{3}/mmc$ (No. 194) Hexagonal 97 77 96 60 0.086 0.545 0.324 4.100
$Fm\bar{3}m$ (No. 225) Cubic 98 96 95 92 0.012 0.033 0.214 0.483
@@
Figs. 7a–g shows the average energy difference to the energy of final structures versus DFT relaxation steps. We neglected the structures whose energy changes and energy change intervals per step during relaxations exceeded 10 eV/atom to eliminate the impact of erroneous steps. One sees that CrystalFormer samples generally reach lower energies in fewer relaxation steps. This is especially true for space groups with higher symmetries. The ability to initialize diverse and high-quality crystal structures enables one to discover more stable materials faster. Fig. 7h shows the histogram of energy above the convex hull constructed by the Materials Project database. The dashed line denotes the criterion $E_{\mathrm{hull}}<0.1$ eV/atom [73] for selecting stable materials. Among these candidates, we found 34 and 12 relaxed structures with CrystalFormer and PyXtal initializations that are not contained in the MP-20 dataset. We summarize them in Table LABEL:table:samples and Table LABEL:table:pyxtal_samples of appendix C.
Table LABEL:table:700-stat lists detailed statistics of structure-relaxed samples in seven representative space groups. Overall, we find that the CrystalFormer generated structures are of higher quality, especially for those space groups with higher symmetry. This observation is supported by the fact that the DFT relaxation often retains the space group symmetry. The root mean squared displacement (RMSD) [16] computed for these converged structures demonstrate CrystalFormer’s superior performance over PyXtal across all 7 space groups. The average energy above the convex hull also confirms the samples generated by CrystalFormer are indeed much closer to the DFT local minimum than PyXtal initialization. CrystalFormer attains superior performance in the high-symmetry space groups compared to the RMSD of 0.11 Å reported in [16] for the MatterGen model trained on the MP-20 dataset with no control on the space-group symmetry.
III.2 Structure-conditioned element substitution
@@
Many crystal structures can be traced back to a few simple, highly symmetrical types. Numerous crystals share the same structural prototype but differ in composition, such as perovskite (ABX3), spinel (AB2X4), fluorite (AX2), and so on. Fig. 8a shows double perovskite crystal structures A2BBX6 which belong to the $Fm\bar{3}m$ (No. 225) space group. There are hundreds of known double perovskites with significant interests in their semiconducting, ferroelectric, thermoelectric, and superconducting properties [77]. Finding more stable materials with this structure prototype using brute force enumeration and high-throughput calculation is a computationally demanding task [78]. We will generate new double perovskites with CrystalFormer and demonstrate its advantage of over standard element substitution methods.
Fig. 8a shows the string representation of double perovskites. To generate candidates of double perovskites, we use CrystalFormer to carry out string in-filling tasks. Since the autoregressive sampling of the atoms is insufficient to take into account non-causal information in the sequence, we employ MCMC to sweep through the sequence and update chemical species and fractional coordinates [79]. The acceptance rate for these MCMC updates makes use of the marginalized probability for elements and fractional coordinates as the lattice parameter that appears at the end of the sequence can be integrated. Only after the MCMC sampling has been thermalized, we sample the lattice parameters autoregressive to account for the adjustment of the unit cell for given atoms and occupations. We use CrystalFormer to generate 100 candidates as the initial DFT relaxation.
@@
[…]
@@
Figure 8: (a) Double perovskites crystal structure. The crystal string representation of double perovskites with blank spaces for chemical elements and the x-coordinate of the atom resides in the ‘e’ position. CrystalFormer generates crystals with double perovskite structures via sequence infilling. (b) The histograms of energy above the convex hull for the relaxed crystal structures. The dashed line indicates the criterion for selecting candidates for the stable materials.
III.3 Plug-and-play materials design
Figure 9: (a) The histogram of band gap and formation energy predicted by MEGNet models for crystal samples generated in the $Fm\bar{3}m$ space group (No. 225). The dashed red lines in the plane indicate target values. The marginals on the side show the shift of the property distributions with respect to the unconditionally generated samples. Note that we scale the 3d histograms for better visualization. (b) The likelihoods of conditioned generated samples compared to the unconditional samples.